What is the focus, directrix, and the equation of the axis of symmetry for y = 1/20x2

Answer by stanbon(75887)

What is the focus, directrix, and the equation of the axis of symmetry
for y = 1/20x2
(x-0)^2 = 20(y-0)
4p = 20
p = 5
---------
Vertex: (0,0)
axis of symmetry: x = 0
-----
directrix: y = -5
------
Focus: (0,5)
------------------
Cheers,
Stan H.
==================